OpenStudy (anonymous):

Sin^2 x + Sin x = 0 Solve the equation for exact solutions over the interval [0, 2Pi]

OpenStudy (anonymous):

solve the equation for exact solutions over the interval [0, 2Pi]

TheSmartOne (thesmartone):

Sin^2 x + cos^2 x = 1 It doesn't equal 0...

OpenStudy (mathstudent55):

Factor out sin x.

TheSmartOne (thesmartone):

Ah the question just got edited lol.

OpenStudy (anonymous):

yeah my bad fellas lol, fixed it

OpenStudy (mathstudent55):

\(\sin^2 x + \sin x = 0\) \(\sin x(\sin x + 1) = 0\) \(\sin x = 0\) or \(\sin x + 1 = 0\) \(\sin x = 0\) or \(\sin x =- 1\)

OpenStudy (mathstudent55):